20:00: we start to prepare for opening the dome...

21:00: PockelsCellsTransmission on LgswUnitSx:

TN Plot Map
('20151023_040836','20151023_040956') 20151023_040836_pcTransmission.jpg 20151023_040836_pcTransmissionMap.jpg
Bad quadrant, we acquire a new dark

22:00: still problems with azimuth drive, we start to acquire LGS at Zentih

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_050438 0, 0, 0 NA NA -1 F F Before acquisition
20151023_052421 0, 0, 0 NA NA -1 F F LGS on PatCams
22:30: azimuth mount drive fixed

22:40: fail again

23:00: we acquire the IM between LM1 and PatCams:

SX LM1 to PatCam IM: 20151023_055300

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_055415 0, 0, 0 NA NA -1 F F
20151023_055501 0, 0, 0 NA NA -1 F F
20151023_060259 0, 0, 0 NA NA -1 F F
LGS are acquired now... We acquire images with LAT to check the geometry on the frame, snapshot is broken...

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_060922 0, 0, 0 NA NA -1 F F
We try a LanFocusStage scan between +68k and -183k steps.

VibComp PupilMirror geometry identification:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_063844 0, 0, 0 NA NA -1 F F VibComp _5_5
20151023_063902 0, 0, 0 NA NA -1 F F VibComp _0_5
20151023_063943 0, 0, 0 NA NA -1 F F VibComp _10_5
20151023_064007 0, 0, 0 NA NA -1 F F VibComp _5_0
20151023_064028 0, 0, 0 NA NA -1 F F VibComp _5_10
23:42: we check LAT pointing model and we also try to calibrate AC to PupilMirror:

SX PM to QC IM: 20151023_072800

00:30: we try to launch beam focus optimization (GOX's procedure):

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_073303 0, 0, 0 NA NA -1 F F
20151023_073319 0, 0, 0 NA NA -1 F F
20151023_073334 0, 0, 0 NA NA -1 F F
20151023_073349 0, 0, 0 NA NA -1 F F
20151023_073404 0, 0, 0 NA NA -1 F F
20151023_073419 0, 0, 0 NA NA -1 F F
20151023_073435 0, 0, 0 NA NA -1 F F
20151023_073450 0, 0, 0 NA NA -1 F F
20151023_073505 0, 0, 0 NA NA -1 F F
best focus is found at -135k steps:

20151023_073303_LFC.png

Then we try the LaserBeamFocusAndCorrector scan (SR's procedure):

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_074106 0, 0, 0 NA NA -1 F F
20151023_074121 0, 0, 0 NA NA -1 F F
20151023_074137 0, 0, 0 NA NA -1 F F
20151023_074151 0, 0, 0 NA NA -1 F F
20151023_074206 0, 0, 0 NA NA -1 F F
20151023_074221 0, 0, 0 NA NA -1 F F
20151023_074237 0, 0, 0 NA NA -1 F F
20151023_074252 0, 0, 0 NA NA -1 F F
20151023_074307 0, 0, 0 NA NA -1 F F
20151023_074322 0, 0, 0 NA NA -1 F F
20151023_074106_LBFCopt.jpg

Best focus @ -160k and corrector @ 350 steps.

Hip9680

1:00: we point a star

2:10: we try again SR beam shaping:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_091212 0, 0, 0 NA NA -1 F F
20151023_091227 0, 0, 0 NA NA -1 F F
20151023_091243 0, 0, 0 NA NA -1 F F
20151023_091259 0, 0, 0 NA NA -1 F F
20151023_091314 0, 0, 0 NA NA -1 F F
20151023_091329 0, 0, 0 NA NA -1 F F
20151023_091344 0, 0, 0 NA NA -1 F F
20151023_091359 0, 0, 0 NA NA -1 F F
20151023_091212_LBFCopt.jpg

Bad measurement, the corrector was @ 350 steps while it should be at 0...
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_091623 0, 0, 0 NA NA -1 F F
20151023_091638 0, 0, 0 NA NA -1 F F
20151023_091653 0, 0, 0 NA NA -1 F F
20151023_091708 0, 0, 0 NA NA -1 F F
20151023_091724 0, 0, 0 NA NA -1 F F
20151023_091739 0, 0, 0 NA NA -1 F F
20151023_091754 0, 0, 0 NA NA -1 F F
20151023_091809 0, 0, 0 NA NA -1 F F
20151023_091824 0, 0, 0 NA NA -1 F F
20151023_091839 0, 0, 0 NA NA -1 F F
20151023_091623_LBFC.png

Same problem: corrector still at 350 steps. We try last time:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_092341 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092358 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092413 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092428 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092444 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092459 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092514 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092530 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092545 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092600 0, 0, 0 NA NA -1 F F LBFC opt
20151023_092600_LBFC.png

5:00: up to now JH tryed to align M3. Pupil wobble went down from 4 subaps to ~1. Back to LGS:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_121050 0, 0, 0 NA NA -1 F F Before ToF scan
20151023_121458 0, 0, 0 NA NA -1 F F ToF adjusted

Another spot optimization:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_122010 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122026 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122041 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122056 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122111 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122126 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122142 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122157 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122212 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122228 0, 0, 0 NA NA -1 F F LBFC opt
20151023_122010_LBFC.png

Then we go back on PatCams, after beam shaping. We close all LGSW internal loops and we acquire Jitter to Pupil mirror IM:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20151023_122608 0, 0, 0 NA NA -1 F F
20151023_130559 0, 0, 0 NA NA 0.788301 T T
20151023_131551 0, 0, 0 NA NA -1 T F
20151023_131638 0, 0, 0 NA NA -1 T F
SX PM to Jitter IM: 20151023_130600 -- MarcoBonaglia - 28 Oct 2015
Topic attachments
I Attachment Action Size Date Who Comment
20151023_040836_pcTransmission.jpgjpg 20151023_040836_pcTransmission.jpg manage 42 K 28 Oct 2015 - 16:56 MarcoBonaglia  
20151023_040836_pcTransmissionMap.jpgjpg 20151023_040836_pcTransmissionMap.jpg manage 38 K 28 Oct 2015 - 16:57 MarcoBonaglia  
20151023_073303_LFC.pngpng 20151023_073303_LFC.png manage 43 K 28 Oct 2015 - 17:11 MarcoBonaglia  
20151023_074106_LBFCopt.jpgjpg 20151023_074106_LBFCopt.jpg manage 44 K 30 Oct 2015 - 15:15 MarcoBonaglia  
20151023_091212_LBFCopt.jpgjpg 20151023_091212_LBFCopt.jpg manage 55 K 30 Oct 2015 - 15:14 MarcoBonaglia  
20151023_091623_LBFC.pngpng 20151023_091623_LBFC.png manage 66 K 28 Oct 2015 - 17:11 MarcoBonaglia  
20151023_092600_LBFC.pngpng 20151023_092600_LBFC.png manage 71 K 28 Oct 2015 - 17:11 MarcoBonaglia  
20151023_122010_LBFC.pngpng 20151023_122010_LBFC.png manage 66 K 28 Oct 2015 - 17:11 MarcoBonaglia  
Topic revision: r2 - 30 Oct 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback